OpenStudy (anonymous):

The area of a rectangle is 21ft , and the length of the rectangle is 1ft less than twice the width. Find the dimensions of the rectangle.

OpenStudy (vishweshshrimali5):

 Area of rectangle = lb

OpenStudy (anonymous):

Try to put this into algebraic expression

OpenStudy (vishweshshrimali5):

You are given that: \[\large{\rm{length} = 2 \times width - 1}\]

OpenStudy (vishweshshrimali5):

Now, we have : \[\large{\rm{length = l; width = b}}\]

OpenStudy (anonymous):

Length: 2x-1 Width: x Area: Length times width, so 2x^2-x=21

OpenStudy (anonymous):

@vishweshshrimali5 double variable is complicated to explain

OpenStudy (vishweshshrimali5):

I was going to write l in terms of b :) Good work @leejaeho2002 :D

OpenStudy (anonymous):

good job both of you @vishweshshrimali5 @leejaeho2002

OpenStudy (vishweshshrimali5):

\[\large{\rm{l = 2b-1}}\] Thus: \[\large{\rm{Area = l\times b = (2b-1)(b)}}\]

OpenStudy (anonymous):

thx

OpenStudy (vishweshshrimali5):

Thanks @study100 :)

OpenStudy (anonymous):

Do u get it homework